# Colorful Minors

### v1

@@ 9.2 From faces to vortices
In this paper we presented some algorithmic consequences of two of our three structural theorems, i.e., theorem 1.5 and theorem 1.6. Clearly, the structure revealed by theorem 1.7 is “intermediate” with respect to that of theorem 1.5 and theorem 1.6. It provides a structural decomposition of -colorful graphs of small bidimensionality, i.e., -colorful graphs excluding a large -colorful rainbow grid as a minor, which we believe should be useful in algorithmic design. As we already mentioned, the algorithmic consequences of theorem 1.7 go beyond the objectives of this paper. However, we outline below some directions on the Steiner Tree problem that are inspired by the current research.
In the introduction, we commented on Steiner Tree as one of the most studied problems on annotated graphs. Using the terminology of this paper, Jansen and Swennenhuis in [45] proved that Steiner Tree, with input a -colorful graph can be solved in time where Is the “best possible” structural parameter on that, when bounded, Steiner Tree can be solved in polynomial time? The answer is negative. This follows by the fact that a polynomial algorithm has been recently given by Groenland, Nederlof, and Koana in [41] for the case where excludes the -colorful rainbow as a colorful minor. As this -colorful graph belongs in the obstruction set because of corollary 6.34, the results of [41] apply to all colorful minors of the -segregated grid on which may be arbitrarily large due to lemma 5.4. We believe that the “limit” is given by the combinatorial bound of theorem 1.7. We conjecture that the bidimensionality of in is the structural frontier for the complexity of Steiner Tree problem in classes that are colorful minor-closed. In particular we conjecture the following complexity dichotomy.
@@
Unless Steiner Tree can be solved in polynomial time in a -colorful minor-closed class if and only if there is some such that excludes the -colorful rainbow -grid as a colorful minor.
Note that the negative part of the above conjecture already follows from the fact that the Steiner Tree problem is NP-complete in planar graphs [34]. We believe that the local structure provided by theorem 1.7 should serve as a starting point for proving the positive part as well. In particular, we now have that all terminals in are confined to the apices and the “few” vortices of the near-embedding of Motivated by the fact that Steiner Tree can be solved in polynomial time when all terminals lie on “few” faces of a planar graph [20, 23, 54], the key challenge becomes how to generalize such face-based techniques to handle apices and vortices in near-embeddings of the kind provided by theorem 1.7. We further anticipate that this methodological approach may lead to algorithmic dichotomies – such as the one conjectured above – for a broader class of problems on annotated graphs, where tractability arises from restricting terminals to few faces of the input graph.
Acknowledgements:
The authors wish to thank Petr A. Golovach, Chun-Hung Liu, and Giannos Stamoulis for fruitful discussions.
@@ Appendix A Some problems where our algorithmic results apply
In this section, we present a selection of natural optimization problems for which our algorithmic results apply. These problems are definable in CMSO and satisfy the conditions required by theorem 8.9 or theorem 8.10.
A.1 Problems where theorem 8.9 can be applied
We now give some examples of optimization problems where theorem 8.9 is applicable. For each of them, there is some formula in CMSO and a choice of such that if the input is the -colorful graph then answer is For all of them it is straightforward to verify that there exists a function such that is a representation of In particular then the question is about paths or cycles, then is enough, while when the problem is defined in terms of some graph then we may pick
@@
A.2 Problems where theorem 8.10 can be applied
We now give two examples of optimization problems where theorem 8.10 is applicable.
@@
In [37] a special case of Ordered Linkability problem was examined, namely the Disjoint Paths Linkability problem that is obtained if is the disjoint union of edges. It was proven in [37] that, when parameterized by this problem is not FPT, unless -linked sets for this particular where introduced by [88] in the late 60s and its graph-theoretical properties have been extensively studied in [62, 47, 79, 8, 21, 50, 83].
As observed in [37], the decision versions of Unordered Linkability and Ordered Linkability are -definable, therefore, from proposition 8.6, they both admit an time algorithm. In other words, while no FPT-algorithms are expected when parameterized by they both become fixed parameter tractable (FPT), when they are further restricted to graphs of bounded Hadwiger number.
According to our results, towards achieving FPT-algorithms, we can do more than bounding the Hadwiger number of the input graph. For this, it can be easily seen that both these problems are folio-representable, therefore, from theorem 1.10, we may deduce that they both admit an time algorithm (recall that for 1-colorful graphs and are the same parameter). This reveals that it is not only the structure of that determines the parameterized complexity of these problems but the way the vertices in are distributed in
@@
As commented in the section 9, by using theorem 8.10 we can construct algorithms running in time for the following decision problem (for every ).
@@
Also, for as commented in section 9.1, we have an algorithm running in time for the following decision problem.
@@
We would like to conclude this section by commenting on how far the simple arguments we explained in section 9.1 can go towards creating constructive algorithms for generalizations of Covering Minor Models of 1-colorful .
Let be any computable colorful minor-monotone 1-colorful graph parameter such that there is a constructive function where, for every 1-colorful graph it holds that Consider now the extension of Covering Minor Models of 1-colorful so that, instead of asking that we ask whether By using the same arguments as we explained in section 9.1, one may construct an FPT-algorithm for this problem.
@@
Appendix B Proof of lemma 7.6
For the proof of lemma 7.6, we use as a starting point a result of Liebenau and Wormald on the number of -regular graphs on vertices. In particular, lemma 7.6 follows by analyzing the asymptotics of the following result.
Proposition B.1 (Liebenau and Wormald [63]).
For all integers and the number of -regular graphs on vertices is asymptotically equal to
Proof of lemma 7.6.
Observe that, by substituting in the formula from proposition B.1, we get
From Stirling’s formula we now get
The two relations above imply that
Notice that the terms and in (4) cancel out and the same holds for all square root terms. By using the fact that (4) implies that
By applying Stirling’s formula again, we may also obtain that
Since the square root terms give and, given that the powers of cancel, the above formula can be rewritten as follows.

### v2

@@ 8.3 From faces to vortices
In this paper we presented some algorithmic consequences of two of our three structural theorems, i.e., theorem˜1.2 and theorem˜1.3. Clearly, the structure revealed by theorem˜1.4 is “intermediate” with respect to that of theorem˜1.2 and theorem˜1.3. It provides a structural decomposition of -colorful graphs of small bidimensionality, i.e., -colorful graphs excluding a large -colorful rainbow grid as a minor, which we believe should be useful in algorithmic design. As we already mentioned, the algorithmic consequences of theorem˜1.4 go beyond the objectives of this paper. However, we outline below some directions on the Steiner Tree problem that are inspired by the current research.
In the introduction, we commented on Steiner Tree as one of the most studied problems on annotated graphs. Using the terminology of this paper, Jansen and Swennenhuis in [46] proved that Steiner Tree, with input a -colorful graph can be solved in time where is the torso treewidth of Is torso treewidth the “best possible” structural parameter on that, when bounded, Steiner Tree can be solved in polynomial time? The answer is negative. This follows by the fact that a polynomial algorithm has been recently given by Groenland, Nederlof, and Koana in [41] for the case where excludes the -colorful rainbow as a colorful minor. As this -colorful graph belongs in the obstruction set because of corollary˜6.34, the results of [41] apply to all colorful minors of the -segregated grid on which torso treewidth may be arbitrarily large due to lemma˜5.4. We believe that the “limit” is given by the combinatorial bound of theorem˜1.4. We conjecture that the bidimensionality of in is the structural frontier for the complexity of Steiner Tree problem in classes that are colorful minor-closed. In particular we conjecture the following complexity dichotomy.
@@
Unless Steiner Tree can be solved in polynomial time in a -colorful minor-closed class if and only if has bounded bidimensionality.
Note that the negative part of the above conjecture already follows from the fact that the Steiner Tree problem is NP-complete in planar graphs [35]. We believe that the local structure provided by theorem˜1.4 should serve as a starting point for proving the positive part as well. In particular, we now have that all terminals in are confined to the apices and the “few” vortices of the near-embedding of Motivated by the fact that Steiner Tree can be solved in polynomial time when all terminals lie on “few” faces of a planar graph [22, 25, 55], the key challenge becomes how to generalize such face-based techniques to handle apices and vortices in near-embeddings of the kind provided by theorem˜1.4. We further anticipate that this methodological approach may lead to algorithmic dichotomies – such as the one conjectured above – for a broader class of problems on annotated graphs, where tractability arises from restricting terminals to few faces of the input graph.
@@ Appendix A Some problems where our algorithmic results apply
In this section, we present a selection of natural optimization problems for which our algorithmic results apply. These problems are definable in CMSO and satisfy the conditions required by theorem˜7.9 or theorem˜7.10.
A.1 Problems where theorem˜7.9 can be applied
We now give some examples of optimization problems where theorem˜7.9 is applicable. For each of them, there is some formula in CMSO and a choice of such that if the input is the -colorful graph then answer is For all of them it is straightforward to verify that there exists a function such that is a representation of In particular then the question is about paths or cycles, then is enough, while when the problem is defined in terms of some graph then we may pick
@@
A.2 Problems where theorem˜7.10 can be applied
We now give two examples of optimization problems where theorem˜7.10 is applicable.
@@
In [37], a special case of Ordered Linkability problem was examined, namely the Disjoint Paths Linkability problem that is obtained if is the disjoint union of edges. It was proven in [37] that, when parameterized by this problem is not FPT, unless -linked sets for this particular where introduced by [90] in the late 60s and its graph-theoretical properties have been extensively studied in [60, 48, 79, 8, 23, 51, 86].
As observed in [37], the decision versions of Unordered Linkability and Ordered Linkability are -definable, therefore, from proposition˜7.6, they both admit an time algorithm. In other words, while no FPT-algorithms are expected when parameterized by they both become fixed parameter tractable (FPT), when they are further restricted to graphs of bounded Hadwiger number.
According to our results, towards achieving FPT-algorithms, we can do more than bounding the Hadwiger number of the input graph. For this, it can be easily seen that both these problems are folio-representable, therefore, from theorem˜1.9, we may deduce that they both admit an time algorithm. This reveals that it is not only the structure of that determines the parameterized complexity of these problems but the way the vertices in are distributed in
@@
As commented in the section˜8, by using theorem˜7.10 we can construct algorithms running in time for the following decision problem (for every ).
@@
Also, for as commented in section˜8.1, we have an algorithm running in time for the following decision problem.
@@
We would like to conclude this section by commenting on how far the simple arguments we explained in section˜8.1 can go towards creating constructive algorithms for generalizations of Covering Minor Models of 1-colorful .
Let be any computable colorful minor-monotone 1-colorful graph parameter such that there is a constructive function where, for every 1-colorful graph it holds that Consider now the extension of Covering Minor Models of 1-colorful so that, instead of asking that we ask whether By using the same arguments as we explained in section˜8.1, one may construct an FPT-algorithm for this problem.